The sum of two numbers is 17. One number is three less than four times the other. What are the numbers?
Answer by solve_for_x(190)

Let x represent one number, and let y represent the other number.

Since the sum of the two numbers is 17, you can write the following
equation:

x + y = 17

Then, since "one number is three less than 4 times the other", you can
write:

x = 4y - 3

Substituting "4y - 3" in place of x in the first equation gives:

4y - 3 + y = 17

5y - 3 = 17

5y = 20

y = 4

Then x = 17 - y = 17 - 4 = 13

Solution: The two numbers are 4 and 13.